OpenStudy (anonymous):

Monthly Sales Goals (in thousands) Stem: Leaf 1 2 5 5 8 2 4 4 6 3 4 6 4 7 8 9 What is the mean monthly sales goal? (And try explain why so I can understand as well?) a) 15000 b)24000 c)25000 d)29000

OpenStudy (amistre64):

each stem represents a "10" spot, each leaf is a "1" spot stem leafs 1 2 5 5 8 represents: 12, 15, 15, 18 or simply, 10+2, 10+5, 10+5, 10+8 for a row value of: 10(4)+(2+5+5+8)

OpenStudy (amistre64):

the mean is still determined as the sum of the parts, divided by how many parts there are

OpenStudy (amistre64):

10(4) + (2+5+5+8) 20(3) + (4+4+6) 30(2) + (4+6) 40(3) + (7+8+9) ^^ there are (4+3+2+3) total parts

OpenStudy (anonymous):

I don't think I really get it fully. I added those four parts and divided and ended up with 87?

OpenStudy (amistre64):

what did you get for the addition of the parts?

OpenStudy (anonymous):

60 74 70 144

OpenStudy (amistre64):

10(4) + (2+5+5+8) 20(3) + (4+4+6) 30(2) + (4+6) 40(3) + (7+8+9) 40 + (18) 60 + (14) 60 + (10) 120 + (24) ----------- 280 + 66 = 346 (4+3+2+3) = 12 my head adding might be a bit off, but i still get pretty close to an option: 346/12

OpenStudy (amistre64):

20, not 18 lol, im 2 off 348/12 = 29

OpenStudy (anonymous):

So it would be d) 29, right

OpenStudy (amistre64):

of course

OpenStudy (anonymous):

cool, i understand it now, thank you so much :)

OpenStudy (amistre64):

good luck